OpenStudy (anonymous):

what is the value of cos(npi/2)

OpenStudy (anonymous):

@harsid Those links don't really relate to the asked problem. Anyway, assuming that \(\bf n\) is an integer, we have the following:\[\bf \cos \left( \frac{ n \pi}{ 2 } \right)=0 \ | \ n=\frac{ \pi }{ 2 }+2 \pi k \ \ or \ \ n=\frac{ 3 \pi }{ 2 }+2 \pi k, \ k \in \mathbb{Z}\]

OpenStudy (anonymous):

You could also say that the values lies within 1 and -1 for real numbers 'n'...